paige is thinking of two numbers that have a product of 180 and difference of 3 what are the two numbers?
Answer by stanbon(75887)

Paige is thinking of two numbers that have a product of 180 and difference of 3 what are the two numbers?
----------------
Solve:
x -y = 3
xy = 180
----
Substitute for "x" and solve for "y":
y(y+3) = 180
----
y^2+3y-180 = 0
(y-12)(y+15) = 0
---
If y = 12, x = 15
If y = -15, y = -12
=========================
Cheers,
Stan H.